OpenStudy (anonymous):

Terrence buys a new car for $20,000. The value of the car depreciates by 15% each year. If f(x) represents the value of the car after x years, which function represents the car’s value?

OpenStudy (anonymous):

f(x)=x/.15, I think

OpenStudy (jack1):

nah, it's annual depreciation, think there'll be a power / exponont in there somewhere one sec

OpenStudy (jack1):

A=P(1−in) A = ? P = 20000 i = interest/depreciation rate n = number of years = x

OpenStudy (jack1):

so \(\Large f(x) = 20000(1-(0.15x))\) ...?